A company is considering a 5-year project to open a new product line. A new machine with an installed cost of $100,000 would be required to manufacture their new product, which is estimated to produce sales of $90,000 in new revenues each year. The cost of goods sold to produce these sales (not including depreciation) is estimated at 43% of sales, and the tax rate at this firm is 39%. If straight-line depreciation is used to calculate annual depreciation, what is the estimated annual operating cash flow from this project each year? (Answer to the nearest dollar.)
Answer #1

Annual depreciation = 100,000 / 5 = 20,000

COGS = 43% of 90,000 = 38,700

Operating cash flow = (90,000 - 38,700 - 20,000)(1 - 0.39) + 20,000

Operating cash flow = 19,093 + 20,000

Operating cash flow = $39,093